# Modelling Cosmic-Ray Transport: Magnetized versus Unmagnetized Motion in Astrophysical Magnetic Turbulence

### v2

@@ 5.3 Generalized Transport Equation
In order to study structure-mediated transport of cosmic rays in astrophysical applications, the physics described by our combined stochastic model need to be encapsulated in a generalized transport equation for the $5+1$-dimensional particle distribution function $f(\boldsymbol{x},\mu,\hat{\omega}_{g},t)$. This transport equation takes the form $\partial_{t}f=\mathcal{D}[f]$, where the generalized transport operator $\mathcal{D}[\cdot]$ can describe a wide range and combination of processes, such as (anomalous) spatial diffusion $\mathcal{D}=D_{xx}\Delta^{(\alpha)}$, pitch-angle diffusion $\mathcal{D}=\partial_{\mu}D_{\mu\mu}\partial_{\mu}$, or momentum diffusion $\mathcal{D}=p^{-2}\partial_{p}p^{2}D_{pp}\partial_{p}$ (with momentum $p\propto\hat{\omega}_{g}$ in our notation) (Metzler & Klafter, 2000; Schlickeiser, 2002). Since transport properties are strong functions of the local plasma conditions, and cosmic rays can exert a dynamically-relevant feedback on the plasma, the transport equation needs to be coupled with the astrophysical simulation for a self-consistent treatment (Bai et al., 2015; Pfrommer et al., 2017; Böss et al., 2023). With such a setup, one could for instance study the effect of streaming on coherent structures (compare, e.g., Rieder & Teyssier, 2017; Sampson et al., 2023).
@@
Possible approaches to construct an operator $\mathcal{D}$, which describes the competition between two modes of transport are contemplated in Section 5.2. We also note that $\mathcal{D}$ must not necessarily be known explicitly, as it can be represented and solved by the corresponding stochastic differential equations (Effenberger et al., 2024; Aerdker et al., 2025). The remaining challenge is then to parametrize our transport model in terms of characteristic turbulence quantities, which may vary across the large-scale simulation domain, such as the turbulence correlation scale, turbulence strength or sonic Mach number.
@@
Such an integrated study can provide data for validating this or other non-standard transport models against observational data. Ultimately, this study should not only be consistent with the data, but provide falsifiable predictions. Observational constraints for intermittent cosmic-ray scattering theories are for instance considered by Butsky et al. (2024) and Kempski et al. (2025).
@@ 6 Summary
To gain a deeper understanding of the transport of highly energetic charged particles, such as cosmic rays, through isotropic structured magnetic turbulence, we studied the motion of test particles in snapshots obtained from a magnetohydrodynamic simulation of a saturated fluctuating dynamo. Based on a careful analysis of the data, we propose a model that separates particle motion into two distinct modes: non-diffusive magnetized transport along strong coherent flux tubes and diffusive unmagnetized transport in weak and highly tangled regions of the magnetic field. We present a stochastic process for each mode, parametrized by separate mean free paths for coherent field line wandering, large-scale mirroring, and unmagnetized scattering. The global transport behaviour emerging from the competition of these processes is described by a combined model, which consists of a stochastic walker that alternates between the two transport modes, where the duration of each segment is sampled from the respective escape-time probability distribution.
The central result of our study is that this combined model accurately reproduces the time-dependent test-particle mean squared displacements, thus providing an explanation for the observed behaviour. Specifically, field-line wandering and large-scale mirroring with long mean free paths along coherent structures give rise to compound subdiffusion, accounting for the pronounced initial ballistic phase and the transient subdiffusive phase. Intermittent scattering and unmagnetized motion with short mean free paths facilitate cross-field transport, giving rise to asymptotically diffusive behaviour.
The regime of particle energies, which is currently numerically accessible, is clearly non-asymptotic, i.e., the energy-dependent scaling of the unconditional mean free path does not yet exhibit the expected low-energy power-law behaviour. However, our results reveal a clear tendency towards the dominance of magnetized motion at small energies, which implies that field-line wandering, non-resonant mirroring, and decoupling of particles from field lines due to intermittent encounters with sharp curvatures are the primary mechanisms for cosmic-ray transport in the asymptotic low-energy regime. We emphasize that their modelling should be done with care to appropriately account for the highly complex nature of structured turbulence, which is evident in the vastly different length scales characterizing individual structures (long coherence lengths and sharp folds), as well as non-trivial correlations between separate structures.
@@ Acknowledgements
We gratefully acknowledge helpful discussions with Martin Lemoine, as well as the warm hospitality of Gaetano Zimbardo and Silvia Perri, who organized stimulating workshops at the University of Calabria in September 2024 and February 2025. J.L. is grateful to Alex Schekochihin and Luca Biferale for the opportunities to present and discuss early stages of this work at the University of Oxford in July 2023 and at the University of Rome Tor Vergata in March 2024, respectively. J.L., P.R. and S.A. would like to acknowledge the Princeton Center for Theoretical Science for hosting the stimulating workshop “Synergistic approaches to particle transport in magnetized turbulence: from the laboratory to astrophysics” in April 2024. The authors also thank the anonymous referee, whose comments helped to improve this manuscript.
@@
[…]
@@
Appendix A Non-dimensionalization of the relativistic Newton-Lorentz equations
@@
with the relativistic particle momentum $\boldsymbol{P}_{t}=\gamma(V_{t})\boldsymbol{V}_{t}$. By writing amplitudes and normalized quantities separately, and re-arranging
@@
where the hats $\hat{\cdot}$ are dropped for notational convenience, and we obtain the normalized gyro frequency
@@
Appendix B Alternative Magnetization Criterion
@@
The results are shown in figure 11, and the decision boundary at $\overline{\!{\delta M}}/\,\overline{\!{M}}\sim 1$ is roughly described by
@@
Although this decision boundary appears to be cleaner when compared to figure 5, transport statistics conditional on $\kappa_{\perp}$ do not yield significant improvements when compared to $\kappa$, especially in respect to the bias of the magnetized data. Additionally, the criterion given by equation (30) does not convey the compelling physical interpretation of $\bar{\kappa}\bar{r}_{g}\sim 1$.
Figure 11: (a) Average of the relative magnetic moment variation $\overline{\!{\delta M}}/\,\overline{\!{M}}$ conditional on the particle gyro radius $\bar{r}_{g}$ and the field-line perpendicular reversal scale $\kappa_{\perp}$. The transition region $\overline{\!{\delta M}}/\,\overline{\!{M}}\sim 1$ between magnetized and unmagnetized transport is approximately described by $\bar{\kappa}_{\perp}\bar{r}_{g}^{1/2}\sim 30$. The joint density $p(\bar{\kappa}_{\perp},\bar{r}_{g})$ is indicated for reference. The remarks for the colour scale of figure 5 apply here as well. (b) Joint density of the relative magnetic moment variation $\overline{\!{\delta M}}/\,\overline{\!{M}}$ and perpendicular magnetization criterion $\frac{1}{30}\bar{\kappa}_{\perp}\bar{r}_{g}^{1/2}$, as well as the conditional average $\left\langle\,\overline{\!{\delta M}}/\,\overline{\!{M}}\middle|\frac{1}{30}\bar{\kappa}_{\perp}\bar{r}_{g}^{1/2}\right\rangle$.
@@ C.1 Langevin equation
with initial conditions $X_{i,0}=0$ and $V_{i,0}\sim\mathcal{N}(0,v_{\mathrm{rms}}^{2}/3)$, and uncorrelated Gaussian noise $\langle\mathop{}\!\mathrm{d}{W}_{i,s}\mathop{}\!\mathrm{d}{W}_{j,s^{\prime}}\rangle=\delta_{ij}\delta(s-s^{\prime})$. The formal solutions of the components of $\boldsymbol{X}_{s}$ and $\boldsymbol{V}_{s}$ can be written as
@@ C.2 Compound subdiffusion
And for the long-time asymptotic behaviour with $\langle X_{s}^{2}\rangle\underset{s\to\infty}{\sim}{2v_{\mathrm{rms}}^{2}}s/{\theta}$ and $\langle s_{t}^{2}\rangle\underset{t\to\infty}{\sim}{2v_{\mathrm{eff}}^{2}t}/{3\theta_{\mu}}$ we can work out
@@
Appendix D Bayesian Optimization
Optimizing the loss function given by equation (23) requires the simulation of a sufficiently large number of samples of our combined stochastic model given by Algorithm 1 at each step of the optimization procedure. The number of samples needs to be large enough in order to produce a sufficiently converged average $\langle\Delta Z_{\mathrm{model},\tau}^{2}\rangle$, but also small enough in order to remain within reasonable computational cost. However, even for a carefully chosen number of samples, evaluating equation (23) remains relatively expensive and noisy.
Figure 12: Landscapes of the loss function equation (23) as estimated by Bayesian optimization, including the expected minimum and estimated confidence intervals.
We therefore turn to Bayesian optimization as implemented in the software package scikit-optimize (Head et al., 2021), which treats evaluations of the loss function as samples drawn from a Gaussian process. Based on previous observations, a cheap acquisition function is optimized, which gives the next point in the parameter space to evaluate. In this way, a reliable estimate of the optimal parameters is achievable for a small number of evaluations of the loss function. The procedure additionally provides a hyper-parameter $\xi$ to balance exploration of uncertain regions of the parameter space against exploitation of regions, which likely contain a minimum of the loss function.
For each $\hat{\omega}_{g}$, we run 64 steps with $\xi=1$ for exploration, followed by 64 steps with $\xi=0.01$ for exploitation of likely minima. For 96 00096\,000 samples of the combined stochastic model, distributed on 96 CPU cores, a single evaluation of the loss function takes about 35 seconds. The resulting loss functions are shown in figure 12, including the expected minima and roughly estimated confidence intervals.
Confidence intervals for Bayesian optimization are not provided by the software package and hard to find in the literature. We thus produce for each $\hat{\omega}_{g}$ a naive estimate based on 480 independent realizations of the underlying trained Gaussian process. The realizations are generated for random points on the parameter grid in the neighbourhood of the expected minimum and the minimum of each realization is recorded. The mean of the minima of the realizations converges to the expected minimum and their standard deviation is taken as confidence bounds on the parameters.

### v3

@@ 5.3 Generalised transport equation
To study structure-mediated transport of cosmic rays in astrophysical applications, the physics described by our combined stochastic model needs to be encapsulated in a generalised transport equation for the $5+1$-dimensional particle distribution function $f(\boldsymbol{x},\mu,\hat{\omega}_{g},t)$. This transport equation takes the form $\partial_{t}f=\mathcal{D}[f]$, where the generalised transport operator $\mathcal{D}[\cdot]$ can describe a wide range and combination of processes, such as (anomalous) spatial diffusion $\mathcal{D}=D_{xx}\Delta^{(\alpha)}$, pitch-angle diffusion $\mathcal{D}=\partial_{\mu}D_{\mu\mu}\partial_{\mu}$, or momentum diffusion $\mathcal{D}=p^{-2}\partial_{p}p^{2}D_{pp}\partial_{p}$ (with momentum $p\propto\hat{\omega}_{g}$ in our notation) (Metzler & Klafter, 2000; Schlickeiser, 2002). Since transport properties are strong functions of the local plasma conditions and cosmic rays can exert a dynamically relevant feedback on the plasma, the transport equation needs to be coupled with the astrophysical simulation for a self-consistent treatment (Bai et al., 2015; Pfrommer et al., 2017; Böss et al., 2023). With such a set-up, one could, for instance, study the effect of streaming on coherent structures (compare, e.g. Rieder & Teyssier, 2017; Sampson et al., 2023).
@@
Possible approaches to construct an operator $\mathcal{D}$, which describes the competition between two modes of transport, are contemplated in Section 5.2. We also note that $\mathcal{D}$ must not necessarily be known explicitly, as it can be represented and solved by the corresponding stochastic differential equations (Effenberger et al., 2024; Aerdker et al., 2025). The remaining challenge is then to parametrise our transport model in terms of characteristic turbulence quantities, which may vary across the large-scale simulation domain, such as the turbulence correlation scale, turbulence strength or sonic Mach number.
@@
Such an integrated study can provide data for validating this or other non-standard transport models against observational data. Ultimately, this study should not only be consistent with the data, but provide falsifiable predictions. Observational constraints for intermittent cosmic-ray scattering theories are, for instance, considered by Butsky et al. (2024) and Kempski et al. (2025).
@@ 6 Summary
To gain a deeper understanding of the transport of highly energetic charged particles, such as cosmic rays, through isotropic structured magnetic turbulence, we studied the motion of test particles in snapshots obtained from a magnetohydrodynamic simulation of a saturated fluctuating dynamo. Based on a careful analysis of the data, we propose a model that separates particle motion into two distinct modes: non-diffusive magnetised transport along strong coherent flux tubes, and diffusive unmagnetised transport in weak and highly tangled regions of the magnetic field. We present a stochastic process for each mode, parametrised by separate mean free paths for coherent field line wandering, large-scale mirroring and unmagnetised scattering. The global transport behaviour emerging from the competition of these processes is described by a combined model, which consists of a stochastic walker that alternates between the two transport modes, where the duration of each segment is sampled from the respective escape-time probability distribution.
The central result of our study is that this combined model accurately reproduces the time-dependent test-particle mean squared displacements, thus providing an explanation for the observed behaviour. Specifically, field-line wandering and large-scale mirroring with long mean free paths along coherent structures give rise to compound subdiffusion, accounting for the pronounced initial ballistic phase and the transient subdiffusive phase. Intermittent scattering and unmagnetised motion with short mean free paths facilitate cross-field transport, giving rise to asymptotically diffusive behaviour.
The regime of particle energies, which is currently numerically accessible, is clearly non-asymptotic, i.e. the energy-dependent scaling of the unconditional mean free path does not yet exhibit the expected low-energy power-law behaviour. However, our results reveal a clear tendency towards the dominance of magnetised motion at small energies, which implies that field-line wandering, non-resonant mirroring and decoupling of particles from field lines due to intermittent encounters with sharp curvatures are the primary mechanisms for cosmic-ray transport in the asymptotic low-energy regime. We emphasise that their modelling should be done with care to appropriately account for the highly complex nature of structured turbulence, which is evident in the vastly different length scales characterising individual structures (long coherence lengths and sharp folds), as well as non-trivial correlations between separate structures.
@@ Acknowledgements
We gratefully acknowledge helpful discussions with Martin Lemoine, as well as the warm hospitality of Gaetano Zimbardo and Silvia Perri, who organised stimulating workshops at the University of Calabria in September 2024 and February 2025. J.L. is grateful to Alex Schekochihin and Luca Biferale for the opportunities to present and discuss early stages of this work at the University of Oxford in July 2023 and at the University of Rome Tor Vergata in March 2024, respectively. J.L., P.R. and S.A. would like to acknowledge the Princeton Center for Theoretical Science for hosting the stimulating workshop “Synergistic approaches to particle transport in magnetized turbulence: from the laboratory to astrophysics” in April 2024. The authors also thank the anonymous referee, whose comments helped to improve this manuscript.
@@
This work was supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) through the Collaborative Research Center SFB1491 “Cosmic Interacting Matters - From Source to Signal” (grant no. 445052434); and the International Space Science Institute (ISSI) in Bern through ISSI International Team project #24-608 “Energetic Particle Transport in Space Plasma Turbulence”. The work of P.R. was initially funded through a Gateway Fellowship and subsequently by the DFG through the Walter Benjamin Fellowship (grant no. 518672034). F.E. acknowledges partial support from NASA LWS grant 80NSSC21K1327. The authors gratefully acknowledge the Gauss Centre for Supercomputing e.V. (www.gauss-centre.eu) for funding this project by providing computing time on the SuperMUC-NG at Leibniz Supercomputing Centre (www.lrz.de) and through the John von Neumann Institute for Computing (NIC) on the GCS Supercomputers JUWELS at Jülich Supercomputing Centre (JSC).
@@
Appendix A Non-dimensionalisation of the relativistic Newton-Lorentz equations
@@
with the relativistic particle momentum $\boldsymbol{P}_{t}=\gamma(V_{t})\boldsymbol{V}_{t}$. By writing amplitudes and normalised quantities separately and re-arranging
@@
where the hats $\hat{\cdot}$ are dropped for notational convenience and we obtain the normalised gyro frequency
@@
Appendix B Alternative magnetisation criterion
@@
The results are shown in figure 11 and the decision boundary at $\overline{\!{\delta M}}/\,\overline{\!{M}}\sim 1$ is roughly described by
@@
Although this decision boundary appears to be cleaner when compared to figure 5, transport statistics conditional on $\kappa_{\perp}$ do not yield significant improvements when compared to $\kappa$, especially in respect to the bias of the magnetised data. Additionally, the criterion given by equation (30) does not convey the compelling physical interpretation of $\bar{\kappa}\bar{r}_{g}\sim 1$.
Figure 11: (a) Average of the relative magnetic moment variation $\overline{\!{\delta M}}/\,\overline{\!{M}}$ conditional on the particle gyro radius $\bar{r}_{g}$ and the field-line perpendicular reversal scale $\kappa_{\perp}$. The transition region $\overline{\!{\delta M}}/\,\overline{\!{M}}\sim 1$ between magnetised and unmagnetised transport is approximately described by $\bar{\kappa}_{\perp}\bar{r}_{g}^{1/2}\sim 30$. The joint density $p(\bar{\kappa}_{\perp},\bar{r}_{g})$ is indicated for reference. The remarks for the colour scale of figure 5 apply here as well. (b) Joint density of the relative magnetic moment variation $\overline{\!{\delta M}}/\,\overline{\!{M}}$ and perpendicular magnetisation criterion $\frac{1}{30}\bar{\kappa}_{\perp}\bar{r}_{g}^{1/2}$, as well as the conditional average $\left\langle\,\overline{\!{\delta M}}/\,\overline{\!{M}}\middle|\frac{1}{30}\bar{\kappa}_{\perp}\bar{r}_{g}^{1/2}\right\rangle$.
@@ C.1 Langevin equation
with initial conditions $X_{i,0}=0$ and $V_{i,0}\sim\mathcal{N}(0,v_{\mathrm{rms}}^{2}/3)$ and uncorrelated Gaussian noise $\langle\mathop{}\!\mathrm{d}{W}_{i,s}\mathop{}\!\mathrm{d}{W}_{j,s^{\prime}}\rangle=\delta_{ij}\delta(s-s^{\prime})$. The formal solutions of the components of $\boldsymbol{X}_{s}$ and $\boldsymbol{V}_{s}$ can be written as
@@ C.2 Compound subdiffusion
Additionally, for the long-time asymptotic behaviour with $\langle X_{s}^{2}\rangle\underset{s\to\infty}{\sim}{2v_{\mathrm{rms}}^{2}}s/{\theta}$ and $\langle s_{t}^{2}\rangle\underset{t\to\infty}{\sim}{2v_{\mathrm{eff}}^{2}t}/{3\theta_{\mu}}$, we can work out
@@
Appendix D Bayesian optimisation
Optimising the loss function given by equation (23) requires the simulation of a sufficiently large number of samples of our combined stochastic model given by Algorithm 1 at each step of the optimisation procedure. The number of samples needs to be large enough to produce a sufficiently converged average $\langle\Delta Z_{\mathrm{model},\tau}^{2}\rangle$, but also small enough to remain within reasonable computational cost. However, even for a carefully chosen number of samples, evaluating equation (23) remains relatively expensive and noisy.
Figure 12: Landscapes of the loss function equation (23) as estimated by Bayesian optimisation, including the expected minimum and estimated confidence intervals.
We therefore turn to Bayesian optimisation as implemented in the software package scikit-optimize (Head et al., 2021), which treats evaluations of the loss function as samples drawn from a Gaussian process. Based on previous observations, a cheap acquisition function is optimised, which gives the next point in the parameter space to evaluate. In this way, a reliable estimate of the optimal parameters is achievable for a small number of evaluations of the loss function. The procedure additionally provides a hyper-parameter $\xi$ to balance exploration of uncertain regions of the parameter space against exploitation of regions, which likely contain a minimum of the loss function.
For each $\hat{\omega}_{g}$, we run 64 steps with $\xi=1$ for exploration, followed by 64 steps with $\xi=0.01$ for exploitation of likely minima. For 96 00096\,000 samples of the combined stochastic model, distributed on 96 CPU cores, a single evaluation of the loss function takes approximately 35 seconds. The resulting loss functions are shown in figure 12, including the expected minima and roughly estimated confidence intervals.
Confidence intervals for Bayesian optimisation are not provided by the software package and hard to find in the literature. We thus produce, for each $\hat{\omega}_{g}$, a naive estimate based on 480 independent realisations of the underlying trained Gaussian process. The realisations are generated for random points on the parameter grid in the neighbourhood of the expected minimum and the minimum of each realisation is recorded. The mean of the minima of the realisations converges to the expected minimum and their standard deviation is taken as confidence bounds on the parameters.